OpenStudy (anonymous):

MEDALS AWARDED!!! Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (anonymous):

is it asking for the dot product?

OpenStudy (anonymous):

yes , dot product of vectors

OpenStudy (anonymous):

ok so just multiply the corresponding components and add them together

OpenStudy (anonymous):

so it would be 2x-2/-8x6,-8x-5?

OpenStudy (anonymous):

(2*-2)+(-8*6)+(-8*-5)=?

OpenStudy (anonymous):

-4+(-42)+40?

OpenStudy (anonymous):

=-4?

OpenStudy (anonymous):

@wilfley ??

OpenStudy (anonymous):

6*8=48 so it would be -4+(-48)+40=-12

OpenStudy (anonymous):

alright, bad multiplication from my part...but is that all?

OpenStudy (anonymous):

that's it, the dot product of vectors is just a number

OpenStudy (anonymous):

i do not use r?

OpenStudy (anonymous):

and one more question...

OpenStudy (anonymous):

Find a ⋅ b. a = <2, 4>, b = <2, 5> = 24??? @wilfley ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you... and one more please lol

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> @wilfley

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

cos(theta)= (dot product of u and v)/(magnitude of u times magnitude of v)

OpenStudy (anonymous):

dot product is 46? how do i find the magnitude?

OpenStudy (anonymous):

@wilfley

OpenStudy (anonymous):

\[\left| u \right|=\sqrt{(6)^2+(-1)^2}\]

OpenStudy (anonymous):

is it 6.08? @wilfley

OpenStudy (anonymous):

46/6.08=7.5?

OpenStudy (anonymous):

you also need the magnitude of v and then multiply that number to the \[\sqrt{37}\] I would give it like that instead of as a decimal unless the question asks for the approximate answer

OpenStudy (anonymous):

then you will have a number that you can use

OpenStudy (anonymous):

\[\frac{ 46 }{ \sqrt{37}\sqrt{65} }=\cos (\theta)\]

OpenStudy (anonymous):

\[\sqrt{47} x \sqrt{65}????\]

OpenStudy (anonymous):

i got .97

OpenStudy (anonymous):

yeah, then take the inverse cosine of that whole mess and you will have your answer

OpenStudy (anonymous):

my answer choices are 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

i got 14 degrees @wilfley

OpenStudy (anonymous):

cool, I get 20.3

OpenStudy (anonymous):

let me show you the work, it will just take a minute

OpenStudy (anonymous):

alright, yea i see what i did wrong... thanks a ton @wilfley

OpenStudy (anonymous):

\[\frac{ 46 }{ \sqrt{37}\sqrt{65} }=\frac{ 46 }{ \sqrt{2405} } so.....\cos^{-1} (\frac{ 46 }{ \sqrt{2405} })=(\theta)\]

OpenStudy (anonymous):

yeaa i see it now... thank you for you help @wilfley !!!!!

OpenStudy (anonymous):

sure!